# A Change Detection Reality Check

### v1

@@ 1 Introduction
The task of change detection from remotely sensed imagery is a canonical and important problem that allows us to analyze how our planet changes over time. It is crucial that change detection methods are highly accurate given their primary applications for building damage assessment disaster response (Gupta et al., 2019a; Sublime & Kalinicheva, 2019) and monitoring of the environment (Cambrin et al., 2024; Watch, 2002). The machine learning and remote sensing fields are currently experiencing their ImageNet (Deng et al., 2009) benchmarking era with new proposed model architectures claiming incremental performance gains on standard benchmark datasets at a rapid rate. For example, the Changer paper (Fang et al., 2023) – first preprint released on Sep. 2022 – is cited by 40 papers, the majority proposing a new architecture for change detection while also not releasing any open-source code or model checkpoints. Many of these proposed methods contain complicated architectural layers and modules specifically designed to better handle the bi-temporal image change detection task (Bai et al., 2023). However, this begs the question if these new methods are actually improvements over generic segmentation architectures or have the benchmarks just been poorly executed? In this paper, we seek to answer this question.
@@ 3.2 Updated WHU-CD Benchmarks
Many papers utilize a preprocessed and randomly split versions of the WHU-CD dataset created in (Bandara & Patel, 2022a) for change detection benchmarks. However these sets have been known to introduce data leakage (Bandara, 2023)$\approx 85\%$ of the test set is included in the train set due to a bug in the preprocessing scripts – which makes them impossible to use to benchmark methods. We perform a new benchmark using the original train and test set splits from the WHU-CD dataset and retrain the BIT, ChangeFormer, and TinyCD models for comparison. We use the same experimental settings described in Section 3.1.

### v2

@@ 1 Introduction
The task of change detection from remotely sensed imagery is a canonical and important problem that allows us to analyze how our planet changes over time. It is crucial that change detection methods are highly accurate given their primary applications for building damage assessment disaster response (Gupta et al., 2019a; Sublime & Kalinicheva, 2019) and monitoring of the environment (Cambrin et al., 2024; Watch, 2002). The machine learning and remote sensing fields are currently experiencing their ImageNet (Deng et al., 2009) benchmarking era with new proposed model architectures claiming incremental performance gains on standard benchmark datasets at a rapid rate. For example, the Changer paper (Fang et al., 2023) – first preprint released on Sep. 2022 – is cited by 56 papers, the majority of them proposing a new architecture for change detection while also not releasing any open-source code or model checkpoints (Ren et al., 2024; Zhang et al., 2024b; Zhou et al., 2024; Li & Wu, ; Liu et al., 2024; Huang et al., 2024; Lin et al., 2024; Wang et al., 2024; Lu et al., 2024; Tan et al., 2024; Xu et al., 2024; Yan et al., 2024; Peng et al., 2023; Wan et al., 2023; Zhao et al., 2023c; Liu et al., 2023c; Quan et al., 2023; Wang et al., 2023; Liu et al., 2023b; Fazry et al., 2023). Many of these proposed methods contain complicated architectural layers and modules specifically designed to better handle the bi-temporal image change detection task (Bai et al., 2023; Zhang et al., 2024a; Yuan et al., 2024; Li et al., 2024; Pang et al., 2024; Zheng et al., 2024; Zhao et al., 2023a; Li et al., 2023; Yan et al., 2023; Chen et al., 2023; Liu et al., 2023a; Zhao et al., 2023b). Notably many of these papers are peer-reviewed and published at reputable conferences and journals. However, this begs the question if these new methods are actually improvements over generic segmentation architectures or have the benchmarks just been poorly executed? In this paper, we seek to answer this question.
@@ 3.2 Updated WHU-CD Benchmarks
Many papers utilize a preprocessed and randomly split version of the WHU-CD dataset created in (Bandara & Patel, 2022a) for change detection benchmarks. However these sets have been known to introduce data leakage (Bandara, 2023)$\approx 85\%$ of the test set is included in the train set due to a bug in the preprocessing scripts – which makes them impossible to use to benchmark methods. We perform a new benchmark using the original train and test set splits from the WHU-CD dataset and retrain the BIT, ChangeFormer, and TinyCD models for comparison. We use the same experimental settings described in Section 3.1.